# Evaluating Task-based Effectiveness of MLLMs on Charts

### v1

@@ Abstract
The ability to automatically interpret and extract insights from charts can support various applications, including assisting individuals with low vision in capturing insights from charts. With advancements in multimodal large language models (MLLMs), particularly GPT-4V, there is potential to meet this demand. However, existing evaluations mainly focus on high-level chart understanding tasks, such as chart captioning, which overlook the low-level data analysis tasks (e.g., characterize distribution) that humans encounter in daily life. In this paper, we explore a forward-thinking question: Is GPT-4V effective at low-level data analysis tasks on charts? To this end, we first curate a large-scale dataset, named ChartInsights, consisting of 89,388 quartets (chart, task, question, answer) and covering 10 widely-used low-level data analysis tasks on 7 chart types. Firstly, we conduct systematic evaluations to understand the capabilities and limitations of 18 advanced MLLMs , which include 12 open-source models and 6 closed-source models. Starting with a standard textual prompt approach, the average accuracy rate across the 18 MLLMs is 36.17%. Among all the models, GPT-4V achieves the highest accuracy, reaching 56.13%. To understand the limitations of multimodal large models in low-level data analysis tasks, we have designed various experiments to conduct an in-depth test of GPT-4V’s capabilities. We further investigate how visual modifications to charts, such as altering visual elements (e.g. changing color schemes) and introducing perturbations (e.g. adding image noise), affect GPT-4V’s performance. Secondly, we present 12 experimental findings. These findings suggest GPT-4V’s potential to revolutionize interaction with charts and uncover the gap between human analytic needs and GPT-4V’s capabilities. Thirdly, we propose a novel textual prompt strategy, named Chain-of-Charts, tailored for low-level analysis tasks, which boosts model performance by 24.36%, resulting in an accuracy of 80.49%. Furthermore, by incorporating a visual prompt strategy that directs GPT-4V’s attention to question-relevant visual elements, we further improve accuracy to 83.83%. Our study not only sheds light on the capabilities and limitations of GPT-4V in low-level data analysis tasks but also offers valuable insights for future research.
@@ 1 Introduction
Figure 1: We evaluate the effectiveness of low-level data analysis tasks on charts in two steps. First, we develop a large-scale dataset tailored for 10 low-level analysis tasks, covering 7 widely-used chart types (top part of the figure). Second, we conduct four main experiments to study the capabilities of GPT-4V on low-level analysis tasks across different scenarios (bottom part of the figure).
@@
Visualization can effectively convey data insights. However, due to the abundance of information a visualization may convey, it is not easy for users to accurately extract the desired information [1, 2]. Therefore, it becomes crucial to automatically help users pinpoint this information [3, 4, 5] based on their needs [6], a process known as chart question answering, or ChartQA for short.
High-Level and Low-Level ChartQA Tasks. ChartQA can generally be categorized into two types: high-level tasks and low-level tasks. High-level tasks typically refer to broader questions such as chart captioning, chart-to-text conversion, etc., while low-level tasks involve more specific inquiries, such as identifying correlations, detecting anomalies, and so forth [6, 3, 7].
@@
Traditionally, ChartQA has been a challenging problem due to the limited capabilities in natural language understanding and the high complexity of chart reasoning. However, recent advancements in multi-modal large language models (MLLMs) have made it possible for users to interact with systems using natural language to extract specific information from data across various modalities. This progress has illuminated new possibilities for ChartQA on different levels of tasks.
Prior Art: High-Level Tasks using MLLMs. Recently, several studies [8, 9, 10, 11, 12, 13, 14] have explored the capabilities of MLLMs in performing high-level ChartQA tasks. The findings reveal that state-of-the-art MLLMs, such as GPT-4V, have demonstrated promising results in addressing high-level tasks. These studies have also outlined a range of intriguing future research directions in this field.
Our Focus: Low-Level Tasks using MLLMs. While the above efforts have concentrated on high-level ChartQA tasks, the effectiveness of MLLMs for low-level data analysis tasks remains underexplored. In this paper, we aim to systematically investigate the capabilities of GPT-4V in addressing 10 low-level data analysis tasks [3, 6]. Our study seeks to answer the following critical questions, shedding light on the potential of MLLMs in performing detailed, granular analyses.
  • Q1: Impact of Textual Prompt Variations.What is the impact of different textual prompts on GPT-4V ’s output accuracy? This question aims to assess the baseline performance and capabilities of GPT-4V in different low-level tasks.
  • Q2: Impact of Visual Variations and Visual Prompts: How do different visual prompts, such as alterations in color schemes, layout configurations (e.g., aspect ratio), and image quality, affect the performance of GPT-4V in low-level tasks?
  • Q3: Impact of Chain-of-Thoughts. Can we enhance basic textual prompts in Q1 with a chain-of-thoughts like approach?
  • Q4: Synergistic Effect of Visual and Textual Prompts: Can the combination of visual and textual prompts lead to enhanced performance in low-level ChartQA tasks with GPT-4V? This question explores the potential for achieving better results by integrating both types of prompts.
Contributions. We make the following notable contribution.
(1) ChartInsights Dataset. We curate a large-scale dataset, ChartInsights for evaluating the low-level data analysis tasks on charts. This dataset features a diverse array of visual variants, visual and textual prompts, and comprehensive metadata, which can be used to investigate the performance of MLLMs in the low-level ChartQA task from different scenarios.
(2) Setting Benchmarks. Our study establishes benchmarks by evaluating GPT-4V, a state-of-the-art MLLM, on 10 low-level ChartQA tasks. This benchmarking effort provides valuable insights into the current capabilities of MLLMs in processing and analyzing chart information.
(3) New Experimental Findings. We conduct a thorough analysis of the experimental results and identify 12 key findings. These insights emphasize the critical role of visual prompts, chart elements, and image quality in successfully performing low-level ChartQA tasks.
(4) Chain-of-Charts. We introduce the Chain-of-Charts strategy, a new textual prompt designed to enhance the reasoning capabilities of GPT-4V in the context of ChartQA. The Chain-of-Charts leverages a series of interconnected question-answer pairs to guide the model.
We also share all our data and code to facilitate further research: https://anonymous.4open.science/r/ChartInsights-D43E
2 Related Work
2.1 Low-Level Analysis Tasks on Charts
Visualization charts offer numerous insights that aid users in performing data analysis tasks. Low-level data analysis tasks typically involve activities requiring direct interpretation and processing of specific visual elements within a chart, such as data retrieval, outlier identification, and correlation determination [7, 3, 15]. Amar et al. [3] identified ten low-level tasks, highlighting the real-world activities users undertake with visualization tools to understand their data. Subsequently, Saket et al. [6] evaluated the effectiveness of five basic charts across ten low-level analysis tasks using two datasets through a crowdsourced experiment. In this paper, we aim to evaluate how effectively GPT-4V can interpret charts by using these ten low-level data analysis tasks as a framework.
2.2 Multimodal Large Language Models
The field of Multimodal Large Language Models (MLLMs) is experiencing rapid advancements, with efforts concentrated on developing artificial intelligence systems capable of processing and producing multi-modal content, including text, images, videos, and more. Early research such as CLIP [16] demonstrated the effective combination of visual and linguistic information through contrastive learning, while subsequent work like DALL-E [17] further showcased the potential of Transformer [18] architecture in generating images that match text descriptions. Building on these foundational successes, the research community has ventured into refining these models for diverse multi-modal applications, employing strategies like fine-tuning and prompt-based learning. For example, VisualGPT [19] and BLIP [20] have been adapted for Visual Question Answering (VQA) tasks, significantly enhancing their multi-modal task performance. Concurrently, the development of various benchmarks [21, 22, 23, 24, 25], including MME [26], has been crucial. These benchmarks provide a wide array of tasks and datasets, facilitating a comprehensive evaluation of MLLMs’ abilities across different contexts. In this paper, we try to harness the off-the-shelfMLLMs for low-level data analysis tasks on charts.
2.3 MLLMs for Chart Question Answering
With the advancements in MLLMs, such as GPT-4V, it becomes increasingly promising to automatically comprehend charts and extract insights according to user queries [1, 2, 4, 27]. This process is known as chart question answering, i.e. ChartQA for short. Recent research efforts have focused on understanding the capabilities of MLLMs in performing ChartQA tasks. These studies can be categorized into two groups: evaluation studies and the construction of datasets for ChartQA.
Evaluating MLLMs on ChartQA Tasks. Several recent studies [8, 9, 10, 11, 12, 13, 14] have attempted to leverage the capabilities of MLLMs to perform high-level ChartQA tasks such as chart captioning and chart-to-text. For example, Huang et al. evaluated the capabilities of representative MLLMs, such as GPT-4V and Bard (i.e. Gemini) [28], on chart captioning tasks. Their findings indicated that GPT-4V faces challenges in generating captions that accurately reflect the factual information presented in charts. Moreover, these studies have highlighted various promising directions for future research in this field. Diverging from the emphasis on high-level tasks in previous works, our research uniquely targets low-level ChartQA tasks [3, 6].
ChartQA Datasets. In the last decade, several ChartQA datasets have been presented [29, 30, 11, 12, 2, 31, 32, 13, 8], as shown in Section 3 Table 1. For example, ChartBench [11] includes 2.1K charts for four types of ChartQA tasks. However, a gap remains evident in the landscape of existing ChartQA datasets: none are tailored to comprehensively evaluate the 10 low-level tasks identified as critical to the ChartQA task. Moreover, to conduct more customized evaluations, such as modifying the visual elements or adding a visual prompt, we need access to the metadata (e.g. the underlying data) of the charts, not just the chart images. Therefore, we curate a large-scale dataset ChartInsights, which consists of a total of 89,388 quartets, each including a chart, a specified task, a corresponding query, and its answer.
3 ChartInsights for Low-level Analysis on Charts
In this section, we will first discuss the design goals for curating datasets for low-level tasks (Section 3.1). We will then provide details of constructing ChartInsights (Section 3.2). We will close this section by elaborating on the characteristics of ChartInsights (Section 3.3).
3.1 Design Goals
G1: Supporting Low-level Data Analysis Tasks. Our first goal is to facilitate the support of 10 low-level data analysis tasks [3, 6]. This focus addresses a critical gap in existing ChartQA datasets, which often overlook the granularity required to fully understand and interact with the data presented in charts.
G2: Evaluating Visual and Textual Variants on Charts. We highlight the critical role of visual variants (e.g. color, size, shape) in data visualizations, which are key to conveying and interpreting information effectively. Despite their importance, these variants are often neglected in existing ChartQA datasets and evaluations. Our goal is to address this issue by incorporating a diverse array of visual variants, including varying chart elements, image quality, and visual prompts. In addition, we also want to investigate the impact of different textual prompts on the low-level analysis task.
G3: Making Metadata Available. The third goal tackles the prevalent issue of inaccessible data and metadata in current ChartQA datasets. By offering comprehensive access to each chart’s metadata, such as the source data, chart type, and visual element specifics (like color schemes and labels), our dataset enhances analytical depth into chart design’s impact on ChartQA performance.
Figure 2: An Overview of ChartInsights
3.2 ChartInsights Construction
To fulfill our three design goals, our construction process begins with the collection of charts with metadata from existing datasets. Next, we meticulously assign specific low-level data analysis tasks to appropriate chart types. Lastly, we develop diverse textual prompt strategies, along with visual variants and prompts, tailored to each chart. Note that we save all metadata during the construction process, which can make the users customize their dataset based on ChartInsights easily.
As shown in Figure 2-(a), the construction of our ChartInsights consists of five steps: Candidate Charts Selection, Low-Level Tasks Generation, Textual Prompts Design, Visual Variants Generation, and Visual Prompts Design.
Step 1: Candidate Charts Selection. In order to more comprehensively evaluate the ability of MLLMs on low-level data analysis tasks, and to conduct more detailed and extended experiments, the datasets (tabular data) and visualization charts we collected need to meet the following three requirements: First, these datasets should contain the original metadata of the chart such as the underlying data for rendering, allowing us to create customized reasoning tasks based on the metadata data. Second, the charts in these datasets should contain data labels, because the lack of data labels will greatly limit the types of low-level tasks. Third, these datasets should contain both simple and complex charts so that the difficulty of the charts is reasonable.
After considering the above three aspects and the characteristics of existing datasets, we chose to extract charts and corresponding metadata from the two existing datasets, namely ChartQA [2] and nvBench [33]. Then, we get a total of 2K high-quality charts as well as their metadata as our initial dataset. The initial dataset contains a total of 7 types of charts, namely stacked bar charts, grouped bar charts, basic bar charts, line charts, grouped line charts, scatterplots, and pie charts.
Step 2: Low-level Tasks Generation. Next, we design a set of low-level tasks for the collected charts. We follow the approach of previous works on designing low-level tasks for charts [7, 3, 6], resulting in 10 low-level tasks in this paper, as shown in top of Figure 1. We group the 10 low-level tasks into three categories, namely Analysis, Search, and Query, based on their purpose and required reasoning abilities [7].
Next, we should decide which tasks are applicable to which types of charts. We will follow the recommendations on the task-based effectiveness of humans to assign the tasks to each chart type [6]. Finally, we have 22,347 (chart, task, question, answer).
Step 3: Textual Prompts Design. In order to better explore the impact of different prompting methods on GPT-4V. We have designed 4 textual prompt methods, namely Fill-in-the-Blank, Multiple-choice, Yes-or-No, and Error Correction prompts. 1) For Fill-in-the-Blank prompt, we maintain the asking method of the initial question and set the answer format for Fill-in-the-Blank prompt; 2) For Multiple-choice prompt, we still maintain the asking method of the initial question, but at this time we will provide a list of choices for GPT-4V, which usually contains one correct answer and two wrong answers, and tells GPT-4V to choose the answer from the options; 3) For Yes-or-No prompt, we first change the initial question to a true or false question and tell GPT-4V whether it needs to be answered correctly or Wrong; and 4) For Error Correction prompt, we put the wrong answer into the original question with a certain probability and change it into a statement.
We expand the above 4 textual prompt variants on the 22,347 (chart, task, question, answer), and thus produce 89,388 (chart, task, question, answer) at the end.
Step 4: Visual Variants Generation. Visual variants (e.g. color, size, shape) of a chart play a key role in delivering insights, but these variants are often overlooked in existing ChartQA datasets and evaluations, and thus we aim to bridge this gap. To this end, we vary the chart elements and add image noise to vary the chart quality.
Figure 4: Three Types of Visual Prompts.
@@ 3.2 ChartInsights Construction
Task Levels Datasets #-Task Types #-Chart Types #-Charts #-Questions #-Questions / #-Charts Metadata Available?
High DVQA [31] 3 1 300K 3.5M 11.7
ChartQA [2] 4 3 4.8K 9.6K 2
Mix FigureQA [30] 6 5 120K 1.5M 12.9
ChartLlama [12] 7 10 11K 160K 14.5
ChartBench [11] 4 11 2.1K 16.8K 8
Low ChartInsights (ours) 10 7 2K 89K 44.5
Step 5: Visual Prompts Design. Kong et al. [34] presented five types of graphical overlays to enhance users’ capabilities in performing data analysis tasks such as extraction and comparison of numerical values. Intuitively, we want to verify whether overlays would have a positive impact on GPT-4V’s performance. Therefore, we design three types of visual prompts (i.e. graphical overlays) for the charts.
Figure 5: The Effectiveness of GPT-4V across 10 Low-level Tasks and 4 Textual Prompts
3.3 ChartInsights Characteristics
The ChartInsights dataset contains 89,388 (chart, task, query, answer) ChartQA samples across 7 chart types for 10 low-level data analysis tasks on charts. Figure 2(b) shows the proportion of 10 low-level tasks (3 task groups) in ChartInsights. Among them, the analysis task group is the largest, accounting for 42.46%. This task group examines the reasoning power of multi-modal large models on charts. Figure 2(c) shows the distribution of 10 low-level tasks and 7 chart types. Note that because we consider the effectiveness of different chart types for different tasks, for some tasks, we did not allocate chart types in Step 2.
Comparison with Existing Datasets. As shown in Table 1, our dataset differs from existing ChartQA datasets [2, 29, 31, 30] in three aspects. First, our dataset covers 10 low-level data analysis tasks across 7 chart types. Second, we have designed a variety of textual prompts, visual variants, and visual prompts for each chart type and task, enabling us to perform numerous ChartQA tasks from different perspectives to evaluate GPT-4V comprehensively. Third, we make all metadata relevant to the charts available, facilitating future research.
@@ 4 Experiments
In this section, we randomly selected 20% of the data from ChartInsights as the test set. The test set contains 17,552 (charts, tasks, questions, answers) samples, involving 400 charts, covering 7 chart types and 10 low-level tasks. The basic textual prompts in the test set is shown in 1 Q1. We first use the test set to evaluate 18 advanced MLLMs and find that GPT-4V has the best performance among all models. The performance of each model on 10 tasks is shown in Table 2. On this basis, we want to better explore the limitations and boundaries of MLLMs in Low-Level Analysis tasks. Taking GPT-4V as a representative research object, in addition to studying the impact of text cue changes, we will also systematically evaluate the performance of GPT-4V on the low-level ChartQA task through three steps: Q2: Impact of visual Prompts, Q3: Impact of Chain-of-Charts, and Q4: Synergistic Effect of Visual and Textual Prompts.
The overall result of GPT-4V is shown in Figure 5, which comprises four heatmap subfigures, labeled (a), (b), (c), and (d), each corresponding to one of the four evaluation methods discussed previously. These heatmaps visualize the performance of GPT-4V on various low-level ChartQA tasks under different prompt conditions. The progression from subfigures (a) to (d) clearly indicates the incremental benefits of incorporating visual prompts, Chain-of-Charts prompts, and their combination, culminating in the most effective approach for improving GPT-4V’s performance in low-level ChartQA tasks.
In the subsequent sections, we will delve into a detailed analysis of each method: Section4.1.2 focuses on the evaluation of GPT-4V using textual prompt variations alone; Section 4.2 explores the impact of visual prompt variations on GPT-4V’s performance; Section 4.3 is dedicated to the evaluation of GPT-4V with Chain-of-Charts prompts; and Section 4.4 covers the combined approach, where both textual and visual prompts are integrated.
4.1 Textual Prompt Evaluation Models on ChartInsights
Experimental Settings As discussed in Section 3.3, we use 17,552 testing samples to evaluate MLLMs. Then, we analyze the answers of MLLMs, compare them with Ground Truth, and calculate the accuracy.
4.1.1 Overall Evaluation on Various MLLMs
In Table 2 and  3, we found that among the 18 models, the performances of closed-source models are far superior to those of open-source models, and the average accuracy rate of these 18 models is calculated to be 38.25%. Among them, VisCPM  [35]has the worst performance at 26.19%, while GPT-4V has the best performance at 56.13%.In all models, GPT-4V achieves its best performance in seven out of 10 tasks. We believe this is because although some open-source models may outperform current cutting-edge proprietary models such as GPT-4V or Gemini-Pro  [28] on certain specific tasks after being fine-tuned on particular datasets, on more general multimodal datasets, closed-source models like GPT-4V still possess strong generalization capabilities and show clear advantages in aspects such as logical reasoning. Finding-1: Closed-source models exhibit far superior generalization performance in low-level analysis tasks compared to open-source models.
Table 2: Accuracy scores of 18 MLLMs across 10 tasks on ChartInsights

### v3

# ChartInsights: Evaluating Multimodal Large Language Models for Low-Level Chart Question Answering

@@ Abstract
Chart question answering (ChartQA) tasks play a critical role in interpreting and extracting insights from visualization charts. While recent advancements in multimodal large language models (MLLMs) like GPT-4o have shown promise in high-level ChartQA tasks, such as chart captioning, their effectiveness in low-level ChartQA tasks (e.g., identifying correlations) remains underexplored. In this paper, we address this gap by evaluating MLLMs on low-level ChartQA using a newly curated dataset, ChartInsights, which consists of 22,347 (chart, task, query, answer) covering 10 data analysis tasks across 7 chart types. We systematically evaluate 19 advanced MLLMs, including 12 open-source and 7 closed-source models. The average accuracy rate across these models is 39.8%, with GPT-4o achieving the highest accuracy at 69.17%. To further explore the limitations of MLLMs in low-level ChartQA, we conduct experiments that alter visual elements of charts (e.g., changing color schemes, adding image noise) to assess their impact on the task effectiveness. Furthermore, we propose a new textual prompt strategy, Chain-of-Charts, tailored for low-level ChartQA tasks, which boosts performance by 14.41%, achieving an accuracy of 83.58%. Finally, incorporating a visual prompt strategy that directs attention to relevant visual elements further improves accuracy to 84.32%.
ChartInsights: Evaluating Multimodal Large Language Models for Low-Level Chart Question Answering
 Yifan Wu1thanks: Equal contribution, Lutao Yan311footnotemark: 1, Leixian Shen2, Yunhai Wang4, Nan Tang1,2, Yuyu Luo1,2thanks: Yuyu Luo is the corresponding author 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3South China University of Technology,  4Renmin University of China ywu012@connect.hkust-gz.edu.cn, yuyuluo@hkust-gz.edu.cn https://chartinsight.github.io/
@@ 1 Introduction
Visualization charts can effectively convey data insights, but the abundance of information they provide makes it challenging for users to efficiently and accurately extract desired information Xie et al. (2024); Li et al. (2024b); Luo et al. (2023). Automated chart question answering (ChartQA) Zeng and Battle (2023); Ye et al. (2024) is crucial to help users pinpoint relevant information based on their intents Amar et al. (2005); Saket et al. (2019).
Figure 1: Examples of Two Types of ChartQA Tasks
High-Level and Low-Level ChartQA Tasks.
ChartQA tasks can generally be categorized into two types: high-level tasks and low-level tasks. High-level tasks focus on questions that require understanding the overall context or summary of the chart, involving broader, goal-oriented inquiries that seek to understand overarching trends or patterns. In contrast, low-level tasks focus on specific, detail-oriented inquiries that seek precise data points or comparisons within the chart, involving straightforward, factual information retrieval Saket et al. (2019); Amar et al. (2005). For example, as shown in Figure 1, given the same line chart showing rainfall in different months of the same year, a high-level task might ask about the cyclical or seasonal trends in the chart, while a low-level task would be more focused on the data itself in the chart, such as asking how much total rainfall there was in July and August.
@@
Traditionally, ChartQA has been a challenging problem due to the limited capabilities in natural language understanding and the high complexity of chart reasoning Li et al. (2024a); Tang et al. (2024); Liu et al. (2024c). Fortunately, recent advancements in multimodal large language models (MLLMs) have made it possible for users to interact with systems using natural language to extract specific information from data across various modalities. This progress has illuminated new possibilities for ChartQA on different levels of tasks.
@@
Task Levels Datasets #-Task Types #-Chart Types #-Charts #-Queries #-Queries/#-Charts Metadata?
High-level ChartQA DVQA Kafle et al. (2018) 3 1 300K 3.5M 11.7
ChartQA Masry et al. (2022) 4 3 4.8K 9.6K 2
Mix FigureQA Kahou et al. (2018) 6 5 120K 1.5M 12.9
ChartLlama Han et al. (2023) 7 10 11K 160K 14.5
ChartBench Xu et al. (2023) 4 11 2.1K 16.8K 8
Low-level ChartQA ChartInsights (ours) 10 7 2K 22K 11.2
Prior Art: High-Level ChartQA with MLLMs.
Recent studies have explored the capabilities of MLLMs in performing high-level ChartQA Cheng et al. (2023); Zhou et al. (2023); Huang et al. (2023); Xu et al. (2023); Han et al. (2023); Xia et al. (2024); Masry et al. (2023). The findings reveal that state-of-the-art MLLMs like GPT-4o have demonstrated promising results in addressing high-level tasks, and have outlined future research directions.
Our Focus: Low-Level ChartQA with MLLMs.
Existing evaluations of ChartQA primarily focus on high-level tasks, such as chart captioning, while overlooking low-level ChartQA tasks (e.g., characterizing distributions) that humans frequently encounter in daily life. Specifically, studies from the visualization and visual analysis community have well-defined 10 widely used low-level ChartQA tasks Amar et al. (2005); Saket et al. (2019). Thus, our study seeks to systematically evaluate the effectiveness of MLLMs in addressing these 10 low-level ChartQA tasks.
Contributions.
The key contributions are:
(1) ChartInsights Dataset. We curate ChartInsights, the first dataset for evaluating low-level data analysis tasks on charts. ChartInsights includes diverse chart variants, textual and visual prompts, and comprehensive metadata, enabling the investigation of MLLMs’ performance across various low-level ChartQA scenarios.
(2) Comprehensive Evaluations. Our study establishes benchmarks by evaluating 19 MLLMs on 10 low-level ChartQA tasks, providing valuable insights into the current capabilities of MLLMs in processing and analyzing chart information.
(3) New Experimental Findings. We summarize 12 experimental findings, highlighting the importance of visual prompts, chart elements, and image quality in performing low-level ChartQA tasks.
(4) Chain-of-Charts. We introduce the Chain-of-Charts strategy, a new textual prompt designed to enhance MLLMs’ reasoning capabilities in ChartQA tasks by leveraging a series of interconnected question-answer pairs to guide the model.
2 ChartInsights Dataset
Figure 2: Low-level Tasks vs. Chart Types
Since no dataset exists for low-level ChartQA tasks, we construct a large-scale dataset, ChartInsights, to systematically evaluate the performance of MLLMs in these tasks. The construction details of ChartInsights are provided in Appendix A.2.
Overview of ChartInsights.
ChartInsights contains 22,347 (chart, task, query, answer) samples across 7 chart types for 10 low-level data analysis tasks on charts Amar et al. (2005). Figure 2 shows the distribution of 10 low-level tasks and 7 chart types. Please refer to Appendix B for more details.
Comparison with Existing Datasets.
As shown in Table 1, ChartInsights differs from existing ChartQA datasets by emphasizing low-level data analysis tasks. It covers 10 distinct tasks across 7 chart types, including 2000 charts with an average of 11.2 queries each. Additionally, all relevant metadata are available, making ChartInsights a valuable resource for future research in ChartQA. Please refer to Appendix A.1 for further discussions.
@@ 3 Experiments
3.1 Experimental Design
Figure 3: An Overview of Experimental Settings
As shown in Figure 3, we utilize our ChartInsights to systematically evaluate the effectiveness of MLLMs in 10 low-level ChartQA from different angles. The experiments are designed as follows:
Exp-1 Benchmarking MLLMs: We start by benchmarking the performance of widely used MLLMs across 10 low-level ChartQA tasks involving 7 different types of charts. This experiment establishes a baseline for understanding the capabilities and limitations of MLLMs in low-level ChartQA tasks.
Exp-2 Impact of Question Types: We analyze how different question types influence MLLMs interactions, helping to identify which types elicit the most accurate and informative responses.
Exp-3 Textual Prompt Strategies: We investigate the effect of various textual prompt strategies, such as Chain-of-Thoughts, on MLLMs performance.
Rethink.
Prior evaluations on high-level ChartQA tasks Cheng et al. (2023); Zhou et al. (2023); Huang et al. (2023); Xu et al. (2023); Han et al. (2023) primarily focused on optimizing textual prompts. Given that ChartQA involves both “reading” and “understanding” chart images, we conduct an in-depth exploration of the impact of visual modification on charts and prompts.
Exp-4 Impact of Visual Prompts: We conduct an in-depth exploration of the impact of visual prompts on MLLMs performance to understand how guiding the MLLMs’ attention to specific visual elements can enhance its analytical capabilities.
Exp-5 Impact of Chart Variations: We vary chart elements to analyze how changes in color schemes, view sizes, and legends affect the performance.
Exp-6 Impact of Image Quality: We evaluate the effect of image quality by introducing various levels of image perturbations, such as noise and resolution changes, to understand the robustness of MLLMs in handling low-quality charts.
Exp-7 Synergistic Effects of Different Strategies: Finally, we explore the synergistic effects of combining different question types, textual prompts, and visual prompts to enhance the overall performance of MLLMs in low-level ChartQA tasks.
Next, we discuss the main results and findings in Section 3.2 and lessons learned in Appendix D.